OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 0) and a radius of 5.

hartnn (hartnn):

Equation of circle with centre h.k and radius r is just \( \large (x-h)^2 + (y-k)^2 = r^2\) here, centre h,k = 0,0 r =5 just plug in values!

OpenStudy (anonymous):

@hartnn could you help with that showing the steps

hartnn (hartnn):

not much steps, you have formula, you have values just plug them in!

OpenStudy (anonymous):

@hartnn not sure how to do that, :( thats why I am asking for help with this problem

hartnn (hartnn):

you are having problems with plugging in values ? or something else ?

OpenStudy (anonymous):

(x−h)2+(y−k)2=r2 (x-0) (y-0) r = 5 @hartnn

hartnn (hartnn):

if you plug h = k = 0, and r = 5 in the formula you get \((x-0)^2 + (y-0)^2 = 5^2\) now just simplify! :)

hartnn (hartnn):

whats 5^2 = ..?

OpenStudy (anonymous):

ok thanks I got it from here @hartnn

hartnn (hartnn):

good! welcome ^_^